# RAILGUN: A Unified Convolutional Policy for Multi-Agent Path Finding Across Different Environments and Tasks

### v1

@@ I Introduction
We present the first centralized learning-based method for MAPF, called RAILGUN, which generates actions based on maps rather than individual agents. The core idea of RAILGUN is to generate a directed graph in which each node has exactly one outgoing edge at every timestep. This design enables our method to handle any number of agents on the map. Additionally, we use a convolutional neural network (CNN) as the model backbone which produces outputs of the same dimensions as the input features. This allows RAILGUN to accommodate maps of varying sizes. In summary, our contributions are as follows:
@@ III-B Lifelong MAPF
Compared to the MAPF problem, Lifelong MAPF (LMAPF) continuously assigns new target locations to agents once they have reached their current targets. In LMAPF, agents do not need to arrive at their targets simultaneously. There are three main approaches to solving LMAPF: solving the problem as a whole [29], using MAPF methods but replanning all paths at each specified timestep [30, 31], and replanning only when agents reach their current targets and are assigned new ones [32, 33]. Some algorithms consider the offline setting in LMAPF, where all tasks are known in advance. Examples include CBSS [34], which applies Traveling Salesman Problem (TSP) methods to plan task orders, and a four-level hierarchical planning algorithm [35] that incorporates MILP and CBS. However, these LMAPF methods also face the same scalability problem as MAPF methods.
@@ III-C Learning-based MAPF
Most other methods focus on the solver side, using imitation learning (IL), reinforcement learning (RL), or both. One early learning-based method for MAPF is PRIMAL [11] which is trained by RL and IL. It is a decentralized algorithm that relies on an FOV around an agent to generate the actions of that agent. MAPF-GPT [15] is a GPT-based model for MAPF problems, trained by IL on a large dataset. Other approaches incorporate communication mechanisms in a decentralized manner, such as GNN [38] and MAGAT [13], which employ Graph Neural Networks (GNNs) for communication, and SCRIMP [14], which uses a global communication mechanism based on transformers.
@@
Figure 1: RAILGUN Inference Overview: On the left side, there is one current state along with all related input features of size $(n,m,1)$. These features are then stacked along the last channel to construct the input feature $F_{in}$ of size $(n,m,k)$. In this example, we have $n=2$, $m=4$, and $k=4$. On the right side, the input feature $F_{in}$ is fed into a CNN-based neural network, which outputs action probabilities $F_{out}$ of size $(n,m,5)$. We sample from $F_{out}$ to obtain actual actions and then apply the corresponding actions to each agent.
Figure 2: This is an example of how an agent-based solution relates to a series of specialized graphs. The upper-left figure illustrates a testcase we aim to solve, along with a graph where green nodes and orange edges represent map connectivity. The other figures show a valid MAPF solution for this testcase, where agent $1$ should yield to agent $2$. At each timestep, each node in the connectivity graph has only one outgoing edge. Here, we draw edges only for nodes occupied by agents, as the outgoing edge for other nodes could be any of the available edges.
@@ IV-A RAILGUN Overview
In this section, we introduce our RAILGUN method. First, we discuss why it is difficult to design a learning algorithm for centralized MAPF where policies are agent-based. When focusing on generating actions based on agent features, we need to provide a neural network with at least the agent’s start location, goal location, and additional features, amounting to $k$ scalar variables ($k\geq 4$) for one agent. Then the total number of features is at least $kN$. Consider that the maximum number of agents could be $N=|V|\approx nm$, where $n$ and $m$ are the 2D map dimensions. If we want to handle all possible numbers of agents on a specific map, the total feature size would be $knm$. This dependence on map size means that we cannot create a policy to cover all different maps if we construct the features agent by agent. That is why there is no centralized learning-based solver and all learning-based MAPF solvers adopt a decentralized approach with a limited FOV for each agent [11, 12, 13, 14, 15].
@@
Our insight is that in a valid MAPF solution, there will be no collision, which means there can be at most one agent in each map grid cell in each timestep. At any timestep, each agent chooses one of the five edges of its grid cell as its action. Therefore, if we remove all edges that the agents do not use at each timestep, we find that a valid MAPF solution can be viewed as a series of specialized graphs. As shown in Figure 2, these specialized graphs have exactly one edge in every occupied grid cell. Once such a directed graph is given, no MAPF solver is needed, as there is only one possible transition at each timestep. The sequence of these specialized graphs then constitutes a valid MAPF solution.
@@
After converting the agent-based solution into a representation as a series of specialized graphs, we use a CNN network to address the challenge of generating these specialized graphs and generalizing across different maps, which we discussed in the previous paragraphs. In this paper, we use standard U-Net architecture [41] for RAILGUN, where the input feature is $F_{in}$ with size $(n,m,k)$ and the output feature is $F_{out}$ with size $(n,m,5)$. Here, $k$ represents the number of feature channels based on the feature design, and $(n,m)$ represents the map size.
@@
As an example shown in Figure 1, to encode an agent’s current location as a feature, we construct a tensor $F_{cur}$ with size $(n,m,1)$. In this tensor, $F_{cur}[i][j]=\text{{idx}}$ if the agent idx is at position $(i,j)$ in the map; otherwise, $F_{cur}[i][j]=0$. Stacking all such feature tensors along the last dimension forms $F_{in}$ with size $(n,m,k)$. $F_{out}[i][j]$ represents the probability distribution over all possible actions at grid cell position $(i,j)$. We use 5 channels because each agent can take one of up to five different actions at each timestep. Thus, if an agent is located at grid cell $(i,j)$, its action probabilities are stored in $F_{out}[i][j]$.
@@ IV-B Model Architecture
We use U-Net as our model backbone mainly due to its success in diffusion models and its ability to output feature maps that are the same size as inputs. We employ the standard U-Net architecture, which consists of a total of five layers. The encoder begins with an initial layer containing 64 channels. At each subsequent layer, the number of channels is doubled while the size of the feature maps is halved. In the decoder, this process is reversed, with the number of channels halved and the spatial resolution doubled at each layer. Notably, bilinear interpolation is not employed in the decoder; instead, we use deconvolution as in original U-Net. At the final layer, the number of channels is reduced to 5, corresponding to the maximum number of possible actions. We should also note that since U-Net uses CNN layers in the encoder, which progressively reduce the spatial dimensions of the feature maps, there is a minimum required input size to ensure valid downsampling operations. For small maps, padding is needed. The resulting RAILGUN model contains approximately 30 million FP32 parameters.
@@ IV-C Feature Selection
As shown in Figure 1, we construct the input features from multiple components. We employ five types of features: the map, current locations, goal locations, cost-to-goal, and gradients of cost-to-goal (the last feature is not shown in Figure 1 due to space constraints). For the map feature, we use 1 to represent non-traversable grid cells and 0 to represent traversable grid cells. For the current and goal locations, we use the agent’s index to indicate which agent occupies a grid cell; otherwise, the grid cell is set to 0. We also attempted encoding agent indices as binary vectors; however, this produces excessively large input features, rendering the model too large to train.
We also use the precomputed shortest path cost as the cost-to-goal feature for each agent which is a widely used feature in learning-based methods, as shown in Figure 1. The gradients of the cost-to-goal, represents the potential direction of next action, are determined by the changes of the cost-to-goal distances. Specifically, we define the changes in cost-to-goal distances from an agent’s current cell $(i,j)$ to its adjacent cells as $\delta_{\text{left}},\delta_{\text{right}},\delta_{\text{up}},\delta_{\text{down}}$. $\delta<0$ indicates that the agent is approaching the goal location; vice versa. The resulting direction, denoted by $\mathbf{g}_{ij}=\left(\Delta x_{ij},\Delta y_{ij}\right)$, consists of horizontal and vertical components. For the horizontal component, $\Delta x_{ij}$ is computed as shown below:
$\Delta x_{ij}=\begin{cases}0&\text{if }\delta_{\text{left}}\geq 0\text{ and }\delta_{\text{right}}\geq 0,\\[4.30554pt] 1&\text{if }\delta_{\text{left}}\geq 0\text{ and }\delta_{\text{right}}<0,\\[4.30554pt] -1&\text{if }\delta_{\text{left}}<0\text{ and }\delta_{\text{right}}\geq 0,\\[4.30554pt] \text{random}(\pm 1)&\text{if }\delta_{\text{left}}<0\text{ and }\delta_{\text{right}}<0,\end{cases}$
@@
Figure 3: Examples of POGEMA-tested maps. The six metrics—Performance, Coordination, Scalability, Cooperation, OOD, and Pathfinding—are evaluated on the following map sets: Maze, Random, Maze, Random, Warehouse, Puzzle, Cities-tiles, and Cities. Note that Cities-tiles are 64×64 areas selected from larger Cities maps with dimensions of 256×256.
@@ V-A Training and Testing Settings
We use POGEMA [16] benchmark to test our method. POGEMA includes several different metrics, allowing a fair multi-fold comparison. For data collection, our training data is primarily generated by LaCAM-v1 [5]. The model is trained with cross-entropy loss and a batch size of 256. We utilize the AdamW optimizer [42] with $\beta$ values set to $(0.9,0.999)$ and a weight decay of $10^{-3}$. The training process achieves convergence in only six hours, leveraging the power of four NVIDIA A100 GPUs.
For training data, we randomly generate 180 maze maps with $32\times 32$ size, each with varying obstacle densities and maze shape. For each map, we randomly generate {2, 5, 20, 40, 60} MAPF scenarios with {16, 32, 64, 96, 128} agents respectively, for a total of 127 scenarios for each map. We use LACAM-v1 [5] to compute the reference paths for all scenarios as the training data.
All experiments [^1] Our video samples are available in the supplementary material. We will also open-source our code after the review process. were conducted on a system running Ubuntu 22.04.1 LTS equipped with an AMD Intel i9-12900K CPU, 128GB RAM and NVIDIA RTX 3080. For the testing phase, the POGEMA benchmark provides a total of 3,376 test cases featuring six different types of maps shown in Figure 3, varying numbers of agents, and different map sizes. POGEMA use six metrics, namely, Performance, Coordination, Scalability, Coopeartion, Out-of-Distribution (OOD) and Pathfinding. The relevant equations are as follows:
@@
[…]
Performance, OOD and Cooperation metrics primarily evaluate solution quality and success rate on different maps. $SoC_{best}$ represents the best SoC performance achieved among all tested algorithms.
@@
Coordination=1# of $\text{Coordination}=1-\frac{\text{\# of collisions}}{\lvert\text{agents}\rvert\times\text{episode\_length}}$
@@ V-B Testing Results
Figure 4 presents the performance metrics for RAILGUN and the baseline methods. The learning-based methods include VDN [43], QPLEX [44], SCRIMP [14], IQL [45], QMIX [46], DCC [47], MAMBA [48], Switcher [49], Follower [50], and MATS-LP [51]. All these baseline methods are decentralized methods. LaCAM-v3 [28] and RHCR [52] serve as the search-based algorithm baselines in MAPF and LMAPF problems.
In Figure 4(a), we observe that RAILGUN achieves high scores across all six metrics and delivers the best performance in four metrics compared to other learning-based methods. RAILGUN attains the highest score in the Scalability metric because it generates specific directed graphs at each timestep, ensuring that runtime depends only on map size rather than the number of agents in theory. However, even though RAILGUN outperforms or matches the scores of other learning-based methods in most areas, it still exhibits a significant gap with LaCAM in SoC-related metrics. This outcome is expected, as RAILGUN is trained on data generated by LaCAM-v1, and LaCAM-v1 is not designed to achieve the best SoC performance. Mimicking LaCAM-v1 is the top priority of RAILGUN rather than producing a valid solution with the lowest SoC.
@@
TABLE I: MAPF Scores on Warehouse: Makespan is the latest agent arrival time. Bold text represents the best score except for LaCAM. “-” represents 0 in CSR and 128 in Makespan.
As shown in Figure 4(b), when testing on LMAPF, a completely zero-shot task for RAILGUN, RAILGUN achieves only moderate scores in throughput-related metrics since none of the training data was optimized for throughput. However, this zero-shot test also demonstrates RAILGUN’s strong generalization ability across different tasks. RAILGUN also attains high scores in Pathfinding, Coordination, and Scalability. These strengths and weaknesses suggest that, although RAILGUN’s overall solution quality remains an issue, it can produce valid solutions in a diverse set of scenarios. Thus, we believe using a dataset optimized for the cost function of interest, combined with applying a task-specific reward function for fine-tuning via RL after the SL process, will help improve the overall solution quality.
@@
Figure 5 presents detailed CSR (see caption) and SoC. Even for unseen maps (Cities-tiles) and larger agent numbers (192 and 256), RAILGUN outperforms other learning-based methods, achieving up to 60% CSR. This also shows RAILGUN’s strong zero-shot generalization ability in new maps and new agent numbers. Furthermore, we observe that DCC attains a better SoC, despite having a lower CSR. This indicates that in DCC, only a few agents fail to reach their goal locations and the path lengths are shorter than those produced by RAILGUN, highlighting that generating valid solutions is a higher priority for RAILGUN.
In Table I, we present the CSR, SoC, and Makespan metrics (see caption) of different algorithms for the Warehouse map. We observe a similar pattern where RAILGUN achieves a higher CSR score but a lower SoC compared to DCC. However, when considering Makespan, RAILGUN outperforms DCC. This indicates that RAILGUN is capable of finding relatively short solutions. As Makespan reflects the latest arrival time, many agents arriving before the last ones contribute to a higher SoC. This may be due to congestion situations, where many agents have a dead lock, and RAILGUN requires additional time to resolve the congestion [^2] We have a demonstration in our video supplementary material.. For LMAPF, as shown in Figure 6, RAILGUN’s throughput score is better than those of VDN, IQL, and MAMBA. Although it does not achieve the best throughput score overall, its scalability is impressive. Figure 6 also demonstrates that as the number of agents increases, the average runtime per agent decreases.

### v2

@@ I Introduction
We present RAILGUN the first centralized learning-based method for MAPF which generates actions based on maps rather than individual agents. The core idea of RAILGUN is to generate a directed graph in which each node has exactly one outgoing edge at every timestep. This design enables our method to handle any number of agents on the map. Additionally, we use U-Net [16] as the model backbone which produces outputs of the same dimensions as the input features. This allows RAILGUN to accommodate maps of varying sizes. In summary, our contributions are as follows:
@@ III-B Lifelong MAPF
Compared to the MAPF problem, Lifelong MAPF (LMAPF) continuously assigns new target locations to agents once they have reached their current targets. In LMAPF, agents do not need to arrive at their targets simultaneously. There are three main approaches to solving LMAPF: solving the problem as a whole [30], using MAPF methods but replanning all paths at each specified timestep [31, 32], and replanning only when agents reach their current targets and are assigned new ones [33, 34]. Some algorithms can solve LMAPF in an offline setting where all tasks are known in advance. include CBSS [35], which applies Traveling Salesman Problem (TSP) methods to plan task orders. However, these LMAPF methods also face the same scalability problem as MAPF methods.
@@ III-C Learning-based MAPF
Most other methods focus on the solver side, using imitation learning (IL), reinforcement learning (RL), or both. Learning-based solves can make end-to-end decisions using all available information and can be trained on various data types (e.g., MAPF, TAPF, LMAPF). In contrast, search-based methods often require multi-stage decomposition with hand-crafted heuristics. One early learning-based method for MAPF is PRIMAL [11] which is trained by RL and IL. It is a decentralized algorithm that relies on an FOV around an agent to generate the actions of that agent. MAPF-GPT [15] is a GPT-based model for MAPF problems, trained by IL on a large dataset. Other approaches incorporate communication mechanisms in a decentralized manner, such as GNN [38] and MAGAT [13], which employ Graph Neural Networks (GNNs) for communication, and SCRIMP [14], which uses a global communication mechanism based on transformers.
@@
Figure 1: RAILGUN Inference Overview: On the left side (features), there is one current state along with all related input features of size $(n,m,1)$. These features are then stacked along the last channel to construct the input feature $F_{in}$ of size $(n,m,k)$. In this example, we have $n=2$, $m=4$, and $k=4$. On the right side (inference), the input feature $F_{in}$ is fed into a CNN-based neural network, which outputs action probabilities $F_{out}$ of size $(n,m,5)$. We sample from $F_{out}$ to obtain actual actions and then apply the corresponding actions to each agent.
Figure 2: This is an example of how an agent-based solution relates to a series of specialized graphs. The upper-left figure illustrates a testcase we aim to solve, along with a graph where green nodes and orange edges represent map connectivity. The other figures show a valid MAPF solution for this testcase, where agent 11 should yield to agent 22. At each timestep, each node in the connectivity graph has only one outgoing edge. Here, we draw edges only for nodes occupied by agents, as the outgoing edge for other nodes could be any of the available edges.
@@ IV-A RAILGUN Overview
In this section, we introduce our RAILGUN method. First, we discuss why it is difficult to design a learning algorithm for centralized MAPF where policies are agent-based. When focusing on generating actions based on agent features, we need to provide a neural network with at least the agent’s start location, goal location, and additional features, amounting to $k$ scalar variables ($k\geq 4$) for one agent. Then the total number of features is at least $kN$. Consider that the maximum number of agents could be $N=|V|\approx nm$, where $n$ and $m$ are the 2D map dimensions. If we want to handle all possible numbers of agents on a specific map, the total feature size would be knmknm. This dependence on map size means that we cannot create a policy to cover all different maps if we construct the features agent by agent. That is why there is no centralized learning-based solver and all learning-based MAPF solvers adopt a decentralized approach with a limited FOV for each agent [11, 12, 13, 14, 15].
@@
Our insight is that in a valid MAPF solution, there will be no collision, which means there can be at most one agent in each map grid cell in each timestep. At any timestep, each agent chooses one of the five edges of its grid cell as its action. Therefore, if we remove all edges that the agents do not use at each timestep, we find that a valid MAPF solution can be viewed as a series of specialized graphs. As shown in Figure˜2, these specialized graphs have exactly one edge in every occupied grid cell. Once such a directed graph is given, no MAPF solver is needed, as there is only one possible transition at each timestep. The sequence of these specialized graphs then constitutes a valid MAPF solution.
@@
After converting agent-based solution into a representation as a series of specialized graphs, we use a CNN network to address the challenge of generating these specialized graphs and generalizing across different maps, which we discussed in the previous paragraphs. The input feature is $F_{in}$ with size $(n,m,k)$, and the output feature is $F_{out}$ with size $(n,m,5)$. Here, $k$ represents the number of feature channels based on the feature design, and $(n,m)$ represents the map size.
@@
As an example shown in Figure˜1, to encode an agent’s current location as a feature, we construct a tensor $F_{cur}$ with size $(n,m,1)$. In this tensor, $F_{cur}[i][j]=\text{{idx}}$ if the agent idx is at position $(i,j)$ in the map; otherwise, $F_{cur}[i][j]=0$. Stacking all such feature tensors along the last dimension forms $F_{in}$ with size $(n,m,k)$. $F_{out}[i][j]$ represents the probability distribution over all possible actions at grid cell position $(i,j)$. We use 5 channels because each agent can take one of up to five different actions at each timestep. Thus, if an agent is located at grid cell $(i,j)$, its action probabilities are stored in $F_{out}[i][j]$. Note that the model only outputs the next action prediction based on the current timestep. Therefore, to obtain a full trajectory, the trained model must be invoked repeatedly.
@@ IV-B Model Architecture
In this paper, we use U-Net for RAILGUN, as it is widely used in diffusion models [41] and includes transposed convolution layers that allows the network to recover the spatial resolution of the input. That’s why RAILGUN can handle input maps of different sizes during inference. We employ the standard U-Net architecture comprising five layers in total. The encoder begins with an initial layer containing 64 channels. At each subsequent layer, the number of channels is doubled while the size of the feature maps is halved. In the decoder, this process is reversed, with the number of channels halved and the spatial resolution doubled at each layer. Notably, bilinear interpolation is not employed in the decoder; instead, we use deconvolution as in original U-Net. At the final layer, the number of channels is reduced to 5, corresponding to the maximum number of possible actions. We should also note that since U-Net uses CNN layers in the encoder, which progressively reduce the spatial dimensions of the feature maps, there is a minimum required input size to ensure valid downsampling operations. For small maps, padding is needed. The resulting RAILGUN model contains approximately 30 million FP32 parameters.
@@ IV-C Feature Selection
As shown in Figure˜1, we construct the input features from multiple components. We employ five types of features: the map, current locations, goal locations, cost-to-goal, and gradients of cost-to-goal (the last feature is not shown in Figure˜1 due to space constraints). For the map feature, we use 1 to represent non-traversable grid cells and 0 to represent traversable grid cells. For the current and goal locations, we use the agent’s index to indicate which agent occupies a grid cell; otherwise, the grid cell is set to 0. We also attempted encoding agent indices as binary vectors; however, this produces excessively large input features, rendering the model too large to train.
We also use the precomputed shortest path cost as the cost-to-goal feature for each agent which is a widely used feature in learning-based methods, as shown in Figure˜1. The gradients of the cost-to-goal, represents the potential direction of next action, are determined by the changes of the cost-to-goal distances. Specifically, we define the changes in cost-to-goal distances from an agent’s current cell $(i,j)$ to its adjacent cells as $\delta_{\text{left}},\delta_{\text{right}},\delta_{\text{up}},\delta_{\text{down}}$. $\delta<0$ indicates that the agent is approaching the goal location. The resulting direction, denoted by $\mathbf{g}_{ij}=\left(\Delta x_{ij},\Delta y_{ij}\right)$, consists of horizontal and vertical components. For the horizontal component, $\Delta x_{ij}$ is computed as shown below:
[…]
@@
Figure 3: Examples of POGEMA-tested maps. The six metrics—Performance, Coordination, Scalability, Cooperation, OOD, and Pathfinding—are evaluated on different map sets. Note that Cities-tiles are 64×64 areas selected from larger Cities maps with dimensions of 256×256.
@@ V-A Training and Testing Settings
We use the POGEMA [17] benchmark to evaluate our method, so for comparison, we only include the methods available in POGEMA. POGEMA includes several different metrics, allowing a fair multi-fold comparison. For data collection, our training data is primarily generated by LaCAM-v1 [5]. The model is trained with cross-entropy loss and a batch size of 256. We utilize the AdamW optimizer [42] with $\beta$ values set to $(0.9,0.999)$ and a weight decay of $10^{-3}$. The training process achieves convergence in only six hours, leveraging the power of four NVIDIA A100 GPUs.
For training data, we randomly generate 180 maze maps with $32\times 32$ size, each with varying obstacle densities and maze shape. For each map, we randomly generate {2, 5, 20, 40, 60} MAPF scenarios with {16, 32, 64, 96, 128} agents respectively, for a total of 127 scenarios for each map. We use LACAM-v1 [5] to compute reference paths for all scenarios as training data, primarily due to its fast data generation speed on large maps.
All experiments [^1] Our code can be found at Github: https://github.com/TachikakaMin were conducted on a system running Ubuntu 22.04.1 LTS equipped with an AMD Intel i9-12900K CPU, 128GB RAM and NVIDIA RTX 3080. For the testing phase, the POGEMA benchmark provides a total of 3,376 test cases featuring six different types of maps shown in Figure˜3, varying numbers of agents, and different map sizes. POGEMA use six metrics, namely, Performance, Coordination, Scalability, Cooperation, Out-of-Distribution (OOD) and Pathfinding. The relevant equations are as follows:
@@
The Performance, OOD, and Cooperation metrics share the same definitions and primarily evaluate solution quality and success rate across different maps. $SoC_{best}$ represents the best SoC performance achieved among all tested algorithms.
@@
$\text{Coordination}=1-\frac{\text{\# of collisions}}{\lvert\text{agents}\rvert\times\text{episode\_length}}$
@@ V-B Testing Results
Figure˜4 presents the performance metrics for RAILGUN and the baseline methods. The learning-based methods include VDN [43], QPLEX [44], SCRIMP [14], IQL [45], QMIX [46], DCC [47], MAMBA [48], Switcher [49], Follower [50], and MATS-LP [51]. All these baseline methods are decentralized methods. LaCAM-v3 [29] and RHCR [31] serve as the search-based algorithm baselines in MAPF and LMAPF problems.
In Figure˜4(a), we observe that RAILGUN achieves high scores across all six metrics. RAILGUN attains the highest score in the Scalability metric because it generates specific directed graphs at each timestep, ensuring that runtime depends only on map size rather than the number of agents in theory. However, even though RAILGUN outperforms or matches the scores of other learning-based methods in most areas, it still exhibits a significant gap with LaCAM in SoC-related metrics. This outcome is expected, as RAILGUN is trained on data generated by LaCAM-v1, and LaCAM-v1 is not designed to achieve the best SoC performance. Mimicking LaCAM-v1 is the top priority of RAILGUN rather than producing a valid solution with the lowest SoC.
@@
TABLE I: MAPF Scores on Warehouse: Makespan is the latest agent arrival time. “-” represents 0 in CSR and 128 in Makespan.
As shown in Figure˜4(b), when testing on LMAPF, a completely zero-shot task for RAILGUN, RAILGUN achieves only moderate scores in throughput-related metrics since none of the training data was optimized for throughput. However, this zero-shot test also demonstrates RAILGUN’s strong generalization ability across different tasks. RAILGUN also attains high scores in Pathfinding, Coordination, and Scalability. These strengths and weaknesses suggest that, although RAILGUN’s overall solution quality remains an issue, it can produce valid solutions in a diverse set of scenarios. Thus, we believe using a dataset optimized for the cost function of interest, combined with applying a task-specific reward function for fine-tuning via RL after the SL process, will help improve the overall solution quality.
@@
Figure˜5 presents detailed CSR (see caption) and SoC. Even for unseen maps (Cities-tiles) and larger agent numbers (192 and 256), RAILGUN outperforms other learning-based methods except SCRIMP, achieving up to 60% CSR. This also shows RAILGUN’s strong zero-shot generalization ability in new maps and new agent numbers. Furthermore, we observe that DCC attains a better SoC, despite having a lower CSR. This indicates that in DCC, only a few agents fail to reach their goal locations and the path lengths are shorter than those produced by RAILGUN, highlighting that generating valid solutions is a higher priority for RAILGUN.
In Table˜I, we present the CSR, SoC, and Makespan metrics (see caption) of different algorithms for the Warehouse map. We observe a similar pattern where RAILGUN achieves a higher CSR score but a lower SoC compared to DCC. However, when considering Makespan, RAILGUN outperforms DCC. This indicates that RAILGUN is capable of finding relatively short solutions. As Makespan reflects the latest arrival time, many agents arriving before the last ones contribute to a higher SoC. This may be due to congestion situations, where many agents have a dead lock, and RAILGUN requires additional time to resolve the congestion [^2] We have a demonstration in our video supplementary material.. SCRIMP achieves the best MAPF performance among learning-based methods, which may indicate that RAILGUN’s supervised learning is not sufficient and that reinforcement learning should also be involved. For LMAPF, as shown in Figure˜6, RAILGUN’s throughput score is better than those of VDN, IQL, and MAMBA. Although it does not achieve the best throughput score overall, its scalability is impressive. Figure˜6 also demonstrates that as the number of agents increases, the average runtime per agent decreases.